• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mitra

    Log entry time 08:03:16 on May 8,2011

    Entry number 358575

    keyword=Owl Shift Summary

    Shift crew:

    SL: Mitra Shabestari
    TO: Chao GU
    3rd: Zinzhan Bai

    Kinematics 6.5: P = 2.845 GeV/c, Theta = 28.0 degrees


    00:00 --- Moved target to Optics
    00:31 --- Moved target to He3
    00:37 --- CODA problem. No start of the run for run# 2474. We had to restart CODA.
    01:08 --- MCC called to let us know that they are having difficulties delivering 120muA current.
    They think that problem comes from injector so we continue data taking with 110muA.
    02:44 --- Beam has become stable after reducing current to 110muA.
    02:49 --- The DAQ rates were very high about 7K and deadtime was also very large.
    We stopped the run. Restarting CODA and deadtime GUI didn't help.
    Called RC; we need to have DAQ expert to solve the problem.
    5:00 --- Alexandre called. He thinks that it can be a module problem. I called MCC
    and asked for escorted access to the hall. Alexandre is on his way.
    05:30 --- Alexandre and Zhihong are done in the hall. It seems that the problem was a loose cable connection
    05:39 --- We got beam back and started taking production data.
    We continued taking production data on He3 till the end of the shift.


    Run list for kinematics 5.6
    =============================

    Run# --- target, current, #of events

    4273 --- Optics, 5muA, 35.5K
    4274 --- He3, 120muA, CODA problem. Run didn't start.
    2475 --- He3, 120-110muA, 967K
    2476 --- He3, 110muA, 1.0M
    2477 --- He3, 110muA, Rates were too high => stopped the run (JUNK RUN)
    2478 --- He3, 110muA, Rates were too high => stopped the run (JUNK RUN)

    2479 - 4294 --- Test and junk runs.

    2495 --- He3, 110muA, 868K
    2496 --- He3, 110muA, 867K
    2497 --- He3, 110muA,... ongoing